OpenStudy (anonymous):

Factor and Show all roots x3 +2x2 −33x=0 I'm pretty ok with factoring but I don't know how to show all roots help would be nice I'm thinking the answer is Factors to: x(x^2 + 2x − 33)= 0 {0,−1+√34, -1 -√34} √= square root

OpenStudy (anonymous):

The roots is what's inside the brackets

OpenStudy (anonymous):

Hi @triciaal I was wondering if maybe you could help me with this it's just that i understood it when you helped me last time but if you can't it's no biggie

OpenStudy (anonymous):

@Cola_161 Your answer is the correct answer. \[x^3+2x^2-33x=0\] \[x\cdot(x^2+2x-33)=0\] \[\boxed{\boxed{x_1=0}}~~or~~(x^2+2x-33)=0\]\[x^2+2x-33=0\]\[\boxed{\boxed{x_2=-1+\sqrt{34}~~and~~x_2=-1-\sqrt{34}}}\]

OpenStudy (anonymous):

Thanks @D3xt3R !

OpenStudy (anonymous):

Nop, could you please vote on "Best Response" ? ;) ^_^